3. Find a formula for the solution of the initial value problem. (a) y + 3y + y = f(t), y(0) = 0, y(0) = 0 (b) y + 4y =Problem C and E please.
Answer #1

3) s) 7+2,14 g = $(4) Auxiliary equation, m72m +150 - ma-1,-1. . General solution , y = (6, + Cat) et + 2 + 2 flt) where D =